Aibell would like to invest $14,000 and have a portfolio expected return of 10.75 percent. She is considering two securities, A and B. A has an expected return of 13.61 percent and B has an expected return of 8.74 percent. How much should she invest in stock A if you invest the balance in stock B?
Answer #1

The return of a portfolio is the weighted average return of the securities which constitute the porfolio

Let weight of Stock A be x.

Stock Weight Expected Return (%) Weight*Expected Return
A x 13.61 13.61x
B 1-x 8.74 8.74*(1-x)

Portfolio Return = \sumWeight*Expected Return

10.75 = 13.61x + 8.74 - 8.74x

13.61x- 8.74x = 10.75-8.74

4.87x = 2.01

x = 2.01/4.87

= 0.41273100616

weight of Stock A = 41.27%

weight of Stock B = 58.73%

Amount invested in Stock A = 14000*41.27%

= 5777.80

Amount invested in Stock B = 14000-Amount invested in Stock A

= 14000-5777.80

= 8222.20